A steadu current of 2.0 A flows in a wire for 16.0s (a) How many coulombsof a charge flow


through the wire? (b) How many electrons flow through the wire for the given time interval?


1
Expert's answer
2022-03-18T09:15:54-0400

(a) How many coulombsof a charge flow through the wire?

"q=It\\
=\rm 2.0\: A\times 16\: s=32\: C"

(b) How many electrons flow through the wire for the given time interval?

"N=\frac{q}{e}\\
=\frac{32}{1.6\times 10^{-19}}=2.0\times 10^{20}"